# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v1

@@ 4.2 Electrification
The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by [Marshall et al.(1995)]:
$E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$ (17)
Charge neutralization processes, occurring after a lightning discharge as described by [Barthe and Pinty(2007)], are governed by the following relationship:
$\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$ (18)
@@
$\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$ (19)
@@
$S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$ (20)
4.3 Atmospheric Background
Our atmospheric background is based on the theory proposed by Hädrich et al. [Hädrich et al.(2020)]. Specifically, we assume that the atmosphere is initially electroneutral, with the charge density, denoted as $\rho$, being zero.
The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture [Anderson(1990)] is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
$\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$ (21)
@@ 4.3 Atmospheric Background
The atmospheric temperature profile [Atmosphere(1975)], $T(h)$, is modeled by a piecewise function that accounts for the lapse rate, including the effect of the inversion layer at a height $h_{1}$. The temperature at a given altitude $h$ is expressed as:
$T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$ (22)
@@
The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation [Houze Jr(2014)], considering the effect of gravity and the ideal gas law. It is given by:
$p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$ (23)
@@
$T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$ (26)
@@
4.4 Electrodynamics
The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations[Ma et al.(1998)], which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$. Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
$\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$ (28)
@@ 4.4 Electrodynamics
To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics. This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$[Kim and Lin(2007)]:
$\nabla^{2}V=-4\pi\rho$ (30)
For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed[Niemeyer et al.(1984)]. The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
$p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$ (31)
Here, $\eta$ is a parameter controlling the spatial concentration of the discharge[Kim and Lin(2007)]. Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation. The sum in the denominator normalizes the probabilities, ensuring that the total probability across all potential locations is unity. This formulation effectively models lightning channel growth, as discharges preferentially propagate along paths with elevated electric potential.
4.5 Fluid Dynamics
The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium [Wendt(2008)]. These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$ (32)
@@ 4.5 Fluid Dynamics
For efficient large-scale atmospheric simulations, we employ the compact Poisson filter method[Rabbani et al.(2022)]. This GPU-friendly solver converts iterative Jacobi steps into a single convolution operation, enabling scalable fluid simulations with controllable accuracy and effective boundary condition handling.
The method solves for a scalar field $\Phi$, which represents the field to be solved in the Poisson equation during the pressure projection.Using the symmetric Poisson kernel $K$, the solution is expressed as:
$K*\Phi=\sum_{i=0}^{r}\mathbf{e}_{i}*_{x}\left(\mathbf{e}_{i}*_{y}\left(\mathbf{e}_{i}*_{z}\Phi\right)\right)$ (33)
Here, $\mathbf{e}_{i}$ are the eigenvectors of $K$, $r$ is the matrix rank, and * denotes convolution along spatial dimensions ($x$, $y$, and $z$).
5 ALGORITHMICS
@@
9:     $u\leftarrow$ Pressure projection using the compact poisson filter Eq.(33)
@@
The theory discussed in the previous section is implemented as a numerical procedure outlined in Algorithm 1. Figure 1 illustrates the interrelationships governing thunderstorm development. We begin by detailing our procedure for modeling the life cycle of a Mesoscale Convective System (MCS). Subsequently, we describe the dynamics of lightning channel growth, which occurs when the electric field threshold is exceeded. Additionally, we present a lightweight parameter toolset designed to facilitate interactive simulation.
5.1 Mesoscale Convective System Cycle
@@
Figure 3: Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns. (a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
@@
Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection. Pressure projection, using the compact Poisson filter (Eq. 33), is applied to ensure stability and enforce flow field divergence constraints. Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs. 1–15).
Boundary conditions for the vapor field ($q_{v}$) include periodic boundaries in the horizontal directions and a Dirichlet condition at the top. For temperature ($\theta$), Dirichlet conditions apply in the horizontal and top directions, with a Neumann condition at the bottom.
@@
5.2 Lightning Channel Growth
The growth of the lightning channel begins at a trigger point, which is randomly selected from locations where the electric field exceeds the threshold. The spatial domain for modeling the lightning channel growth is typically represented at a subgrid scale within the thunderstorm model, using a coarser grid size. [Mansell et al.(2002)]This is because the lightning channel is much more sparse compared to the broader thundercloud structure, and it only appears in a limited region of the sky.
Figure 4: Initial charge configurations for different types of lightning. (a) CG Lightning, (b) IC Lightning. Red regions represent areas where the electric potential is equal to 1.
Drawing inspiration from the DBM model proposed by [Kim and Lin(2007)], the lightning channel growth is terminated once it reaches a positive charge object. This approach provides greater flexibility in controlling the final structure of the lightning channel, while also improving computational efficiency.
For cloud-to-ground (CG) lightning, a positive charge configuration is assumed to exist on the ground, with the electric potential set to one. In contrast, for intracloud (IC) lightning, a radial positive charge distribution is established within the boundary of the lightning solving grid,as shown in Figure 4.
The evolution of the lightning channel is modeled by solving the Poisson equation (Eq. 30), using the Gauss-Seidel Successive Overrelaxation (SOR) method. The solution is performed on a grid, and then the results are converted from grid points to geometrical points. To achieve a more natural branching structure, we apply a dithering technique to the points.
Additionally, we employ a custom edge transport method to differentiate point groupings, allowing the model to separately control the widths of return strokes and dart leaders. Finally, this process yields the lightning geometry for the current simulation frame, which is then prepared for rendering.
6 VISUALIZATION
Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration. The hardware setup includes NVIDIA® GeForce® RTX 3060 Ti GPU,13th Gen Intel® Core™ i5-13600KF processor, and 32GB of RAM.
The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events. As shown in Table LABEL:tab:lightweight_params, we summarize the lightweight parameter set used for the scenes presented in this section. All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
6.1 Thunderstorms Variation
@@
  • Single Cell: This is the simplest form of thunderstorm, typically short-lived and consisting of a single updraft and downdraft cycle. It often forms in isolated conditions and usually dissipates within an hour.
  • Multicell: Multicell thunderstorms are composed of multiple convective cells at various stages of their life cycle. These storms exhibit greater longevity and intensity compared to single cells, as new cells continuously form along the gust front of the system.
  • Squall Line: A squall line is a collection of thunderstorms that align into a long, narrow band, often stretching hundreds of miles. This type of storm is typically associated with intense winds, heavy rainfall, and the characteristic ”shelf cloud” feature that marks the leading edge of the gust front.
  • Supercell: The supercell is the most intense and organized type of thunderstorm, characterized by a rotating updraft known as a mesocyclone. These storms are capable of producing severe weather phenomena, including large hail, damaging winds, and tornadoes.
These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS. The visual comparison of these thunderstorm types is presented in Figure LABEL:fig:thundercloud_comparison, highlighting their structural differences and unique features.
6.2 Severe Weather Phenomena
@@
These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather. The corresponding visualizations are presented in Figure LABEL:fig:weather_simulation.
7 VALIDATION
We follow the evaluation methods in [Shen et al.(2020)] and [Cesana et al.(2019)], tracking altitude-based quantitative data to analyze cloud behavior across atmospheric layers. Results are shown in Figure LABEL:fig:cloud_fraction.
Following the methodology in [Herrera et al.(2021)], we analyze the temporal evolution of four storm types: single cell, multicell, squall line, and supercell. Cloud coverage evolution reveals dynamic distribution changes, while temperature evolution highlights thermal dynamics and energy exchanges.
Following the evaultaion method in [Formenton et al.(2013)],We analyze the lightning flash rate evolution to capture thunderstorm electrical activity. The flash rate rises during storm development and falls to zero as it dissipates, reflecting the storm lifecycle and charge dynamics.
We specifically selected four regions frequently studied in thunderstorms meteorology [Uman(2001), Harris and Carvalho(2018)] to validate our model by comparing its simulation results with real-world weather data. Our focus includes cloud cover dynamics, temperature trends, and relative humidity variations.
Figure LABEL:fig:cloud_real illustrates this comparison across Florida, New Mexico, Japan, and California, showcasing the model’s capability to capture diverse climatic behaviors and atmospheric dynamics.
8 CONCLUSION AND FUTURE WORK

### v3

@@ 4.2. Electrification
The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by (Marshall et al., 1995):
(17) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$
Charge neutralization processes, occurring after a lightning discharge as described by (Barthe and Pinty, 2007), are governed by the following relationship:
(18) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$
@@
(19) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$
@@
(20) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$
4.3. Atmospheric Background
Our atmospheric background is based on the theory proposed by Hädrich et al. (Hädrich et al., 2020). Specifically, we assume that the atmosphere is initially electroneutral, with the charge density, denoted as $\rho$, being zero.
The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture (Anderson, 1990) is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
(21) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$
@@ 4.3. Atmospheric Background
The atmospheric temperature profile (Atmosphere, 1975), $T(h)$, is modeled by a piecewise function that accounts for the lapse rate, including the effect of the inversion layer at a height $h_{1}$. The temperature at a given altitude $h$ is expressed as:
(22) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$
@@
The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation (Houze Jr, 2014), considering the effect of gravity and the ideal gas law. It is given by:
(23) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$
@@
(26) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$
@@
4.4. Electrodynamics
The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations(Ma et al., 1998), which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$. Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
(28) $\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$
@@ 4.4. Electrodynamics
To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics. This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
(30) $\nabla^{2}V=-4\pi\rho$
For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed(Niemeyer et al., 1984). The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
(31) $p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$
Here, $\eta$ is a parameter controlling the spatial concentration of the discharge(Kim and Lin, 2007). Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation.
4.5. Fluid Dynamics
The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium (Wendt, 2008). These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
(32) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$
@@ 4.5. Fluid Dynamics
5. ALGORITHMICS
@@
9:     $u\leftarrow$ Pressure projection
@@
The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1. Figure 3 visualizes the key interrelationships governing thunderstorm development. We first present our methodology for simulating the life cycle of a Mesoscale Convective System (MCS), followed by an explanation of the dynamics of lightning channel growth triggered when the electric field exceeds a defined threshold.
5.1. Mesoscale Convective System Cycle
@@
Figure 5. Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns. (a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
@@
Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection. Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported through advection based on the evolving velocity field. For efficient pressure projection, we employ the compact Poisson filter scheme (Rabbani et al., 2022), a GPU-friendly method for solving large-scale sparse linear systems. Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs. 1–15).
@@
5.2. Lightning Channel Growth
The growth of the lightning channel begins at a trigger point, randomly selected from locations where the electric field exceeds a predefined threshold. Inspired by the DBM model proposed by (Kim and Lin, 2007), the lightning channel stops growing once it connects to a positively charged object. This approach not only enhances flexibility in controlling the final structure of the lightning channel but also improves computational efficiency. The positive charge is distributed along the top of the height field.
The solution is computed on a volumetric grid, and the results are subsequently converted from grid points to geometric points. To create a more natural branching structure, a random dithering technique is applied to these points. The points are then connected to form the final geometry of the lightning branches.
6. VISUALIZATION
Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration. The hardware setup includes NVIDIA® GeForce® RTX A6000 GPU,13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events. As shown in Table 1, we summarize the lightweight parameter set used for the scenes presented in this section. All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
Table 1. Overview of the lightweight parameter set used in the scenes presented in this section. $D$: World dimensions (km); $R$: Volume resolution; $T_{\text{mcs}}$: Computation time in seconds per frame(s).
Fig. Scene $D$ (km) $R$ $T_{mcs}$ (s)
6(a) Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
6(b) Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
6(c) Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
6(d) Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
7(a) Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
7(b) New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
7(c) Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
7(d) California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
6.1. Thunderstorms Variation
@@
  • Single Cell: Single cell thunderstorms, often referred to as “popcorn” convection, are small, isolated storms characterized by their brief lifespan, typically lasting less than an hour. These storms are visually compact with a single updraft and downdraft cycle, forming as isolated towering cumulus clouds.
  • Multicell: Multicell thunderstorms consist of clusters of individual convective cells, each at varying stages of development. They appear as a dynamic structure where new cells continuously form along the gust front, sustaining the system for several hours.
  • Squall Line: A squall line is a linear arrangement of thunderstorms, visually identifiable by a long and narrow band of cumulonimbus clouds. These storms can extend hundreds of miles in length but are typically narrow, often around 10-20 miles wide.
  • Supercell: The supercell is the most visually striking and organized thunderstorm type, dominated by a large, rotating updraft known as a mesocyclone. These storms feature massive, tilted, and rotating cloud structures, often rising up to 50,000 feet. The mesocyclone can span up to 10 miles in diameter and persists for hours.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Refer to caption(e)
Refer to caption(f)
Refer to caption(g)
Refer to caption(h)
Figure 6. Comparison between simulation results (a-d) and realistic observations (e-h) for four storm types: single cell, multicell, squall line, and supercell. This layout highlights the structural similarities between our simulated storms and their real-world counterparts.
These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS. The visual comparison of these thunderstorm types is presented in Figure 6, highlighting their structural differences and unique features.
6.2. Severe Weather Phenomena
@@
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 7. Simulation cloudscapes inspired by real-world weather events. (a) Florida Biscayne National Park, (b) New Mexico Chiricahua Mountains, (c) Japan Fuji-Hakone-Izu National Park, (d) California coastal atmospheric river. These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.
  • Biscayne National Park, Florida (25.3692, -80.243): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
  • Chiricahua Mountains, New Mexico (35.7943, -106.443): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
  • Fuji-Hakone-Izu National Park, Japan (37.745, -119.54): On January 10, 2011, a sudden winter ;p‘thunderstorm formed in this region, resulting in brief but intense snowfall.
  • Monterey Bay, California (36.6197, -121.906): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
@@
These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather. The corresponding visualizations are presented in Figure 7.
7. VALIDATION
We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across different atmospheric layers. The results are presented in Figure 8, which illustrates the relationship between cloud fraction and storm height. Notably, the maximum cloud fraction corresponds to the storm’s peak altitude, consistent with the classical vertical distribution of cumulonimbus clouds in meteorology.
Figure 8. Comparative visualization of cloud fractions under different storm systems. From left to right: single-cell, multi-cell, squall line, and supercell. Each panel highlights the cloud fraction structure typical of these systems.
For further validation, we specifically selected four regions frequently studied in thunderstorm meteorology (Uman, 2001; Harris and Carvalho, 2018). Following the methodology from (Herrera et al., 2021), we examine the temporal evolution of the thunderstorm described in section 6.2. The evolution of cloud coverage is compared with real-world weather data from public sources, showing dynamic changes in the distribution of cloud cover. This comparison helps assess the accuracy of the model in capturing cloud coverage patterns observed in nature.we present the maximum storm development height along with the corresponding cloud fraction, providing a detailed overview of the storm’s vertical structure. This information is used to validate the model’s ability to simulate the typical altitude and distribution of clouds during storm development.
In line with the evaluation method in (Formenton et al., 2013), we analyze the evolution of lightning flash rates to capture the electrical activity of thunderstorms. The flash rate increases as the storm develops and gradually declines to zero as the storm dissipates. This trend reflects the typical lifecycle of a thunderstorm and its charge dynamics. Additionally, the frequency of lightning occurrences is examined in relation to altitude.
Figure 9 provides a comparison across Florida, New Mexico, Japan, and California, illustrating the model’s ability to capture a wide range of climatic behaviors and atmospheric dynamics, and confirming its consistency with observed storm structures and lightning patterns.
8. CONCLUSION AND FUTURE WORK